# Mixing, segregation, and collapse transitions of interacting copolymer rings

### v2

@@ 1 Introduction
We use several techniques to investigate this phase diagram. In section 3, we use rigorous arguments to establish the existence of a phase boundary as $\beta_{m}$ is increased, for all $\beta_{c}<\infty$. We write this phase boundary as $\beta_{m}=\beta_{m}^{*}(\beta_{c})$. In the same section, we show that, if a collapse transition exists as $\beta_{c}$ is increased with $\beta_{m}=0$ (say at $\beta_{c}=\beta_{c}^{*}$), then a transition occurs at the same value of $\beta_{c}$ for all $\beta_{m}<\beta_{m}(\beta_{c}^{*})$. It is thought that the transition at $\beta_{c}^{*}$ is second order and there are good numerical estimates of its location [3, 28]. The estimates are
$\beta_{c}^{*}=\begin{cases}0.2782\pm 0.0070,&\hbox{for collapsing polygons};\\ 0.2779\pm 0.0041,&\hbox{for collapsing walks}.\end{cases}$ (1)
@@ 2 The model
The pair of self-avoiding lattice polygons have total length $2n$, and each has length $n$, both embedded in the cubic lattice, $\mathcal{Z}^{3}$. The polygons are also mutually avoiding and each has a root vertex placed such that the two roots are one lattice unit distance apart. We label the two polygons A and B and this setup can either describe two polymer rings linked together by a single chemical bond or a diblock copolymer with a figure eight connectivity.
@@ 2.1 Preliminary considerations
Above a critical value $\beta_{m}^{*}$, the system is expected to be in a Mixed (M) phase where the two rings are significantly intermingled (the density of mutual contacts $\langle k_{m}\rangle/2n$ is now positive in the thermodynamic limit). The value of $\beta_{m}^{*}$ is $0.31$ for $\beta_{c}=0$  [12] but, for $\beta_{c}\neq 0$, we expect $\beta_{m}^{*}=\beta_{m}^{*}(\beta_{c})$, at least for $\beta_{c}$ sufficiently large.
Finally we note that, to some extent, there is a similarity with the adsorption and collapse transition problem [34, 36, 35]. In particular, the fact that the density of mutual contacts is negligible below the mixing transition could suggest that some features of the boundaries between the phases are similar. For instance, we might guess that the phase boundary between the SE and the SC phases ($\theta$-point) does not depend on the value of $\beta_{m}$ when $\beta_{m}<\beta_{m}^{*}$.
@@ 3 Some rigorous results
Let $p_{n}(k_{c})$ be the number of lattice polygons of length $n$ steps in the $d$-dimensional hypercubic lattice with $k_{c}$ nearest neighbour contacts between vertices. This defines a model of lattice polygons with nearest-neighbour self-interactions and partition function
@@
$\lim_{n\to\infty}\frac{1}{n}\log P_{n}(\beta_{c})=\kappa_{d}(\beta_{c})$ (4)
exists in the hypercubic lattice for all $\beta_{c}<\infty$ [28].
@@
We do not know that $\kappa_{d}(\beta_{c})$ is not analytic, rigorously, at a collapse critical point $\beta_{c}^{*}$. However, if a collapse transition exists at $\beta_{c}=\beta_{c}^{*}$, then $\kappa_{3}(\beta_{c})$ is non-analytic at $\beta_{c}^{*}$. It then follows from Theorem 1 that $\phi(\beta_{m},\beta_{c})$ is also not analytic at $\beta_{c}^{*}$ for all $\beta_{m}\leq\beta_{m}^{*}(\beta_{c}^{*})$. This shows that, if the collapse transition exists, then the phase boundary separating the SE and SC phases in figure 2 must be a straight line. It is not known that the phase boundary separating the SE and M phases is a straight line.
@@ 4 Monte Carlo method
Figure 4: Crossing the Mixed-SC boundary either by keeping fixed $\beta_{c}>\beta_{c}^{*}$ and (a) varying $\beta_{m}$ or by (b) keeping fixed $\beta_{m}>\beta_{m}^{*}$ and varying $\beta_{c}$.
The stationary distribution of the MCC implementation is the product of the Boltzmann distributions along each chain (see reference [29, 28]). The main difficulty in the practical implementation of the MMC scheme is the choice of the sequence of parameter pairs $(\beta_{c}{(j)},)\beta_{m}{(j)}),j=1,2,\ldots,M$. This is because the contiguous pairs $(\beta_{c}{(j)},)\beta_{m}{(j)})$ and $(\beta_{c}{(j+1)},)\beta_{m}{(j+1)})$ should be sufficiently close to allow a non-negligible number of swaps between them to occur. In this study, we will estimate the metric and thermodynamic properties of the model by either fixing $\beta_{m}$ and varying $\beta_{c}$ or fixing $\beta_{c}$ and letting $\beta_{m}$ vary. More specifically, we choose a set of values of $\beta_{m}$ at fixed $\beta_{c}$ or a set of values of $\beta_{c}$ at fixed $\beta_{m}$ which samples the relevant region of the phase diagram and crosses the expected phase boundaries previously described.
For most cases we used $M\approx 25$ parallel chains, and we were able to obtain sufficiently uncorrelated samples for systems of total size $2n\in\{96,200,296,400,600,800\}$. We counted a single iteration as $O(1)$ pivot moves, and $O(n)$ local Verdier-Stockmayer style moves. By spacing the reading of data along each Markov chain, we were able to sample approximately $2.5\times 10^{4}$ conformations that are essentially uncorrelated at each fixed value of the pair $(\beta_{c}^{(j)},\beta_{m}^{(j)})$ for a total of at least $1.25\times 10^{6}$ data points for each value of $n$. The simulations were expensive in terms of CPU time. For example, for $2n=800$ and $M\approx 25$, a total of 25,00025,\!000 uncorrelated conformations were sampled over two months of CPU time on an x86_64x86\_64-based workstation. In figure 3 we show some examples of conformations taken along the chains for different values of the pair $(\beta_{c}^{(j)},\beta_{m}^{(j)})$.
@@ 5.1 The SE-SC phase boundary
Consider first the case that $\beta_{m}=0$. In this event the only nearest-neighbour self-attraction between monomers is when $\beta_{c}>0$. In the event that $\beta_{c}\leq 0$ the model is in a phase of segregated and expanded polygons. Increasing $\beta_{c}>0$ the model is expected to undergo a collapse transition at the $\theta$-point at a critical value $\beta_{c}^{*}>0$. In the case of cubic lattice polygons $\beta_{c}^{*}\approx 0.278$ [28, 3], and one expects this to be also the location of the $\theta$-point in our model when $\beta_{m}\leq 0$.
@@
Figure 6: (a) The locations of the peaks of $Var(k_{c})/2n$ for three values of the mutual attraction parameter $\beta_{m}=0.0,0.20,0.30$. (b) The corresponding locations of the crossings between the curve $\langle R_{g}^{2}\rangle/n^{2/3}$ vs $\beta_{c}$ at $n=400$ and the curves for $n=48,100,148,200$ and $300$. The three values of $\beta_{m}$ are the same as in (a). In both cases the extrapolation is against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ refers to the estimate of $\beta_{c}^{*}$ reported for the standard $\theta$-point in [28].
In figure 6(a) we plot the estimated locations of the maxima in the variances $Var(k_{c})/2n$. This is then extrapolated against $1/\sqrt{n}$ (at the $\theta$-point the finite size crossover exponent $\phi$ is expected to have its mean field value $1/2$). The crossings between the scaled mean square radius of gyration curves $\langle R_{g}^{2}\rangle/n^{2/3}$ for $n=400$ with the corresponding curves estimated for $n=48,100,148,200$ and $300$ are plotted in figure 6(b). The thick segment at $1/\sqrt{n}=0$ in the two graphs marks the estimated location of the cubic lattice $\theta$-point $\beta_{c}^{*}=0.2782\pm 0.007$ in reference [28]. Observe that (i) the $n\to\infty$ extrapolations of the data in both graphs are very close to previous estimates of the cubic lattice $\theta$ transition; and (ii) the phase boundary between the SE and SC phases seems to be independent of the strength of the mutual attraction $\beta_{m}$, at least in the range $\beta_{m}\in[0,\beta_{m}^{*}]\approx[0,0.31]$.
@@ 5.2 SE-M phase boundary
If $\beta_{c}=0$ then the model reduces to two polygons with a mutual nearest neighbour attraction between vertices in different polygons. In this case the model goes through a segregated-mixed phase transition at $\beta_{m}^{*}\approx 0.31$ [12]. The average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, is shown as a function of $\beta_{m}$ (for $\beta_{c}=0$ and $0.28$) in figure 7 for $n=48,100,148,200,300$ and $400$.
@@
The intersecting variance curves in the second row of graphs are consistent with a critical point $\beta_{m}^{*}$ corresponding to a transition into a mixed phase. The intersections are also consistent with the estimate $\beta^{*}_{m}\approx 0.31$.
If $\beta_{m}<\beta_{m}^{*}(\beta_{c}=0)$ then $\langle k_{m}\rangle/2n$ tends to zero as $n$ increases (this is seen in the left-most panel in the first row in figure 7). This behaviour persists also for $0<\beta_{c}<\beta_{c}^{*}$ (as seen in the right-most panel in the top row of figure 7). These results are consistent with two polygons being segregated in space if $\beta_{m}<\beta_{m}^{*}$. Notice that above a particular value of $\beta_{m}$ the curves $\langle k_{m}\rangle/2n$ increase with $n$ in the panels in the top row of figure 7. This increase in mutual contacts characterizes a mixed phase where the two components share the same region in space (when the model transitions as $\beta_{m}$ increases beyond the critical point $\beta_{m}^{*}\approx 0.31$ in figure 2 separating the SE- and M-phases).
@@
Figure 8: (a) The locations of the crossings of the variance per unit length curves $Var(k_{m})/2n$ as a function of $n=400$ with the corresponding curves for $m=48,148,200,300$, (b) As in the left panel but for the curves $d_{cm}$ against $\beta_{m}$. The three values of $\beta_{c}$ are $0.0,0.20$ and $0.28$ (see legend). In both panels data have been extrapolated against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ is the estimate of $\beta_{m}^{*}$ (see reference [12]) for the $\beta_{c}=0$ case.
@@
Figure 9: (a1-a3) Plots of the average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, as a function of $\beta_{m}$ for $\beta_{c}=0.4,0.5$ and $0.75$. These values are well above the theta transition $\beta_{c}^{*}\approx 0.28$ (see the right panel of figure 4). Different symbols refer to different system sizes (see the legend). (b1-b3) Plots of the variance per unit length, $Var(k_{m})/2n$, as a function of $\beta_{m}$ for the same fixed values $\beta_{c}=0.4$, $0.5$ and $0.75$. (c1-c3) Corresponding $\beta_{m}$ dependence of the distance between the centre of mass of the two polygons, $d_{cm}$.
@@
Figure 10: Left panel: The locations of crossings $\beta_{m}^{*}(Var(k_{m})^{cr})$ between the curve $Var(k_{m})/2n$ for $n=400$ (see the middle row of figure 9) and $Var(k_{m})/2n$ for $n\in\{48,100,148,200,300\}$ for values of $\beta_{c}<\beta_{c}^{*}$ (below the $\theta$-point; these are the empty symbols) with those above $\beta_{c}^{*}$ (solid symbols). The locations are plotted as a function of $1/\sqrt{n}$. Notice that the data below the $\theta$-point tends to line up and extrapolate as $n\to\infty$ to the marked solid vertical bar at $1/\sqrt{n}=0$ (this is the estimate for $\beta_{m}^{*}$ in reference [12] for the case $\beta_{c}=0$. In contrast, the data above the $\theta$-point extrapolate to larger values on the vertical axis. Right panel: This is similar to the left panel, but now for the curves $d_{cm}$ in the middle row of figure 9. The values of $\beta_{c}$ are reported in the legend.
@@
The onset of a thermodynamic phase transition between the SE- and M-phases is also suggested by the $\beta_{m}$ dependence of $Var(k_{m})/2n$ in the curves in the middle row of panels in figure 7. Notice the onset of peaks whose heights increase with $n$. Moreover, the curves cross one another in a narrow region of the mutual interaction parameter $\beta_{m}$. This is consistent with the theorem that, for $\beta_{c}<\infty$, and $\beta_{m}<\beta_{m}^{*}$, $\lim_{n\to\infty}Var(k_{m})/2n=0$; see section 3 and theorem 1. That result shows the presence of an asymmetric phase transition when $\beta_{c}<\infty$ somewhat similar to the adsorption transition of a self-avoiding walk anchored to an impenetrable attractive wall [10]. In these cases it is known that the intersections of the curves $Var(k_{m})/2n$ plotted against $\beta_{m}$ provide a faithful finite-size estimate of the location of the phase boundary [16].
@@
A further indication of an SE-M phase boundary is seen in the behaviour of the distance between the centres of mass of the two components, $d_{cm}$, as a function of $\beta_{m}$ (these data are plotted in the bottom row panels in figure 7). In the SE phase, for $\beta_{c}=0$, the hard-core repulsion between vertices in the two polygons induces an entropic repulsive force between the polygons, segregating them when $\beta_{m}<\beta_{m}^{*}$. Here the natural length scale is $n^{\nu}$, where $\nu$ is the metric exponent of the self-avoiding walk. This shows that $d_{cm}\sim n^{\nu}$ in the SE-phase. In the M-phase the two-component polygons share the same region as the attractive forces have overcome the entropic repulsion due to hard core nearest neighour interactions. In this case $d_{cm}$ should approach a constant, or grow slowly with $n$ at a rate slower than $O(n^{\nu})$. These expectations are supported by the data in the left-most panel in the bottom row of figure 7, and the corresponding data for $\beta_{c}=0.28$ seen in the other panel show consistent behaviour with $\beta_{c}=0$, also indicating that the phase boundary between the SE- and M-phases in figure 2 is a straight line.
@@
To determine the approximate location of the critical point $\beta_{m}^{*}$ the crossings between the curves between $Var(k_{m})/2n$ and $d_{cm}$ were determined and plotted against $1/\sqrt{n}$ in figure 8 (assuming the mean field value $1/2$ for the finite size crossover exponent $\phi$ and plotting the intersections between the variance for $n=400$ and for $n=48,100,148,200$ and $300$). Extrapolating these data points indicates a limiting value close to $0.31$ (the thick segment at $1/\sqrt{n}=0$ marks the estimate of the expanded-mixed phase transition $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [12]). A similar extrapolation is reported for the location of the crossings of the curve $d_{cm}(n=400)$ with the corresponding curves obtained for $n=48,100,148,200$ and $300$ in the last row of panels in figure 8. The limiting estimate shown by the thick segment at $1/\sqrt{n}=0$ is again the expanded-mixed phase transition point $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [12]. Similarly to the SE-SC phase boundary, we note that: (i) the $n\to\infty$ extrapolation of the points of both measures is very close to $\beta_{m}^{*}$, and (ii) the SE-M phase transition appears to be independent of the strength $\beta_{c}$ of the self attraction between vertices in the polygons, at least when $0\leq\beta_{c}<\beta_{c}^{*}$.
@@
Figure 12: (a) Comparison of the locations of the maxima of the variance curves $Var(k_{c})/2n$ for various values of $\beta_{m}$. These data extrapolate to $\beta_{m}^{*}$ for data below the M-SE transition (this corresponds to the empty symbols), but to larger values above the M-SE transition (solid symbols) the M-SE transition. (b) Similar to (a), not now for the separation between the centre-of-mass data $d_{cm}$ plotted for various values of $\beta_{m}$. The values of $\beta_{m}$ are reported explicitly in the legend. In both panels data have been plotted as a function of $1/\sqrt{n}$. The vertical bars at $1/\sqrt{n}=0$ refers to the estimates of $\beta_{m}^{*}$ provided in reference [12] for the $\beta_{m}=0$ case.

### v3

@@ 1 Introduction
We use several techniques to investigate this phase diagram. In section 3, we use rigorous arguments to establish the existence of a phase boundary as $\beta_{m}$ is increased, for all $\beta_{c}<\infty$. We write this phase boundary as $\beta_{m}^{*}(\beta_{c})$ in the phase diagram. Note that $\beta_{m}^{*}(\beta_{c})$ is a function of $\beta_{c}$ and is the phase boundary separating the mixed $M$-phase from the segregated $SE$- and $SC$-phases (see figure 2). Similarly, we denote a phase boundary seperating the expanded $SE$-phase and the mixed $M$-phase from the collapsed $SC$-phase by $\beta_{c}^{*}(\beta_{m})$.
If $\beta_{c}=0$ then the model has a transition from the $SE$-phase to the $M$-phase at the critical point $\beta_{m}^{*}(0)$, where we understand that $\beta_{m}^{*}(0)$ is the critical point for the transition when $\beta_{c}=0$. We similarly define the critical point $\beta_{c}^{0}$ separating the $SE$- and $SC$-phases when $\beta_{m}=0$. It is thought that the transition at $\beta_{c}^{0}$ is second order and there are good numerical estimates of its location [3, 27]. The estimates are
$\beta_{c}^{0}=\begin{cases}0.2782\pm 0.0070,&\hbox{for collapsing polygons};\\ 0.2779\pm 0.0041,&\hbox{for collapsing walks}.\end{cases}$ (1)
@@ 2 The model
The pair of self-avoiding lattice polygons has total length $2n$, and each has length $n$, both embedded in the cubic lattice, $\mathcal{Z}^{3}$. The polygons are also mutually avoiding and each has a root vertex placed such that the two roots are one lattice unit distance apart. We label the two polygons A and B and this setup can either describe two polymer rings linked together by a single chemical bond or a diblock copolymer with a figure eight connectivity.
@@ 2.1 Preliminary considerations
Above a critical value $\beta_{m}^{*}(\beta_{c})$, the system is expected to be in a Mixed (M) phase where the two rings are significantly intermingled (the density of mutual contacts $\langle k_{m}\rangle/2n$ is now positive in the thermodynamic limit). The SESE-$M$ transition occurs at the critical point $\beta_{m}^{*}(0)\approx 0.31$ for $\beta_{c}=0$  [12] but, for $\beta_{c}\neq 0$, we expect $\beta_{m}^{*}(\beta_{c})$ to be an increasing function of $\beta_{c}$ for large $\beta_{c}$.
Finally we note that, to some extent, there is a similarity with the adsorption and collapse transition problem [33, 34]. In particular, the fact that the density of mutual contacts is negligible below the mixing transition could suggest that some features of the boundaries between the phases are similar. For instance, we might guess that the phase boundary between the SE and the SC phases ($\theta$-point) does not depend on the value of $\beta_{m}$ when $\beta_{m}<\beta_{m}^{*}(0)$.
@@ 3 Some rigorous results
Let $p_{n}(k_{c})$ be the number of lattice polygons of length $n$ steps in the cubic lattice with $k_{c}$ nearest neighbour contacts between vertices. This defines a model of lattice polygons with nearest-neighbour self-interactions and partition function
@@
$\lim_{n\to\infty}\frac{1}{n}\log P_{n}(\beta_{c})=\kappa_{3}(\beta_{c})$ (4)
exists in the cubic lattice for all $\beta_{c}<\infty$. More generally, this limit also exists in the hypercubic lattice [27], in which case the thermodynamic limit is denoted $\kappa_{d}(\beta_{c})$. In the square lattice, it is denoted $\kappa_{2}(\beta_{c})$. Clearly, $\kappa_{2}(\beta_{c})\leq\kappa_{3}(\beta_{c})$.
@@
We do not know that $\kappa_{d}(\beta_{c})$ is not analytic, rigorously, at a collapse critical point $\beta_{c}^{0}$. However, if a collapse transition exists at $\beta_{c}=\beta_{c}^{0}$, then $\kappa_{3}(\beta_{c})$ is non-analytic at $\beta_{c}^{0}$. It then follows from Theorem 1 that $\phi(\beta_{m},\beta_{c})$ is also not analytic at $\beta_{c}^{0}$ for all $\beta_{m}\leq\beta_{m}^{*}(\beta_{c}^{0})$. This shows that, if the collapse transition exists, then the phase boundary separating the SE and SC phases in figure 2 must be a straight line. It is not known that the phase boundary separating the SE and M phases is a straight line.
@@ 4 Monte Carlo method
Figure 4: Crossing the M-SC boundary by keeping fixed $\beta_{m}>\beta_{m}^{*}(0)$ and varying $\beta_{c}$ (left) or by keeping fixed $\beta_{c}>\beta_{c}^{0}$ and varying $\beta_{m}$ (right).
The stationary distribution of the MCC implementation is the product of the Boltzmann distributions along each chain (see reference [28, 27]). The main difficulty in the practical implementation of the MMC scheme is the choice of the sequence of parameter pairs $(\beta_{c}{(j)},)\beta_{m}{(j)}),j=1,2,\ldots,M$. This is because the contiguous pairs $(\beta_{c}{(j)},)\beta_{m}{(j)})$ and $(\beta_{c}{(j+1)},)\beta_{m}{(j+1)})$ should be sufficiently close to allow a non-negligible number of swaps between them to occur. Swaps between chains were proposed at a rate proportional to the system size ($2n$ – that is the total combined length of the pair of polygons, for $n=48,100,148,200,300,400$).
We proceeded by collecting data of the metric and thermodynamic properties of the model along the chains by either fixing $\beta_{m}$ and varying $\beta_{c}$ or fixing $\beta_{c}$ and letting $\beta_{m}$ vary. More specifically, we choose a set of values of $\beta_{m}$ at fixed $\beta_{c}$ or a set of values of $\beta_{c}$ at fixed $\beta_{m}$ which samples the relevant region of the phase diagram and crosses the expected phase boundaries previously described.
For most cases we used $M\approx 25$ parallel chains, and we were able to obtain sufficiently uncorrelated samples for systems of total size $2n\in\{96,200,296,400,600,800\}$. We counted a single iteration as $O(1)$ pivot moves, and $O(n)$ local Verdier-Stockmayer style moves. By spacing the reading of data along each Markov chain, we were able to sample approximately $2.5\times 10^{4}$ conformations that are essentially uncorrelated at each fixed value of the pair $(\beta_{c}^{(j)},\beta_{m}^{(j)})$ for a total of at least $1.25\times 10^{6}$ data points for each value of $n$. The simulations were expensive in terms of CPU time. For example, for $2n=800$ (each polygon in the pair has length $400$) and $M\approx 25$, a total of 25,00025,\!000 uncorrelated conformations were sampled over two months of CPU time on an x86_64x86\_64-based workstation. In figure 3 we show some examples of conformations taken along the chains for different values of the pair $(\beta_{c}^{(j)},\beta_{m}^{(j)})$.
@@ 5.1 The SE-SC phase boundary
Consider first the case that $\beta_{m}=0$. In this event the only nearest-neighbour self-attraction between monomers is when $\beta_{c}>0$. In the event that $\beta_{c}\leq 0$ the model is in a phase of segregated and expanded polygons. Increasing $\beta_{c}>0$ the model is expected to undergo a collapse transition at the $\theta$-point at a critical value $\beta_{c}^{0}>0$. In the case of cubic lattice polygons $\beta_{c}^{0}\approx 0.278$ [27, 3], and one expects this to be also the location of the $\theta$-point in our model when $\beta_{m}\leq 0$.
@@
Figure 6: Estimating the location of the critical point $\beta_{c}^{0}$ from the locations of peaks in $Var(k_{c})$ (denoted by $\beta_{c}^{*}(Var(k_{c}))$) and the locations of crossing points of mean square radius of gyration curves (denoted by $\beta_{c}^{*}(R_{g})$) – see figure 5 panels (b1) and (b2), and panels (d1) and (d2). (a) The locations of the peaks of $Var(k_{c})/2n$ for three values of the mutual attraction parameter $\beta_{m}=0.0,0.20,0.30$. (b) The corresponding locations of the crossings between the curve $\langle R_{g}^{2}\rangle/n^{2/3}$ vs $\beta_{c}$ at $n=400$ and the curves for $n=48,100,148,200$ and $300$. The three values of $\beta_{m}$ are the same as in (a). In both cases the data are plotted against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ refers to the estimate of $\beta_{c}^{0}$ reported for the standard $\theta$-point in [27].
In figure 6(a) we plot the estimated locations of the maxima in the variances $Var(k_{c})/2n$. These maxima were estimated using a fourth order spline fit to the data points. We plot these maxima against $1/\sqrt{n}$ (at the $\theta$-point the finite size crossover exponent $\phi$ is expected to have its mean field value $1/2$). The crossings between the scaled mean square radius of gyration curves $\langle R_{g}^{2}\rangle/n^{2/3}$ for $n=400$ with the corresponding curves estimated for $n=48,100,148,200$ and $300$ are plotted in figure 6(b). The thick segment at $1/\sqrt{n}=0$ in the two graphs marks the estimated location of the cubic lattice $\theta$-point $\beta_{c}^{0}=0.2782\pm 0.007$ in reference [27]. Observe that (i) the $n\to\infty$ extrapolations of the data in both graphs are very close to previous estimates of the cubic lattice $\theta$ transition; and (ii) the phase boundary between the SE and SC phases seems to be independent of the strength of the mutual attraction $\beta_{m}$, at least in the range $\beta_{m}\in[0,\beta_{m}^{*}(0)]\approx[0,0.31]$.
@@ 5.2 SE-M phase boundary
If $\beta_{c}=0$ then the model reduces to two polygons with a mutual nearest neighbour attraction between vertices in different polygons. In this case the model goes through a segregated-mixed phase transition at $\beta_{m}^{*}(0)\approx 0.31$ [12]. The average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, is shown as a function of $\beta_{m}$ (for $\beta_{c}=0$ and $0.28$) in figure 7 for $n=48,100,148,200,300$ and $400$.
@@
The intersecting variance curves in the second row of graphs are consistent with a critical point corresponding to a transition into a mixed phase. The intersections are also consistent with the estimate $\beta^{*}_{m}(\beta_{c})\approx 0.31$ for $\beta_{c}\in[0,\beta_{c}^{0}]$.
If $\beta_{m}<\beta_{m}^{*}(0)$ then $\langle k_{m}\rangle/2n$ tends to zero as $n$ increases (this is seen in the left-most panel in the first row in figure 7). This behaviour persists also for $0<\beta_{c}<\beta_{c}^{0}$ (as seen in the right-most panel in the top row of figure 7). These results are consistent with two polygons being segregated in space if $\beta_{m}<\beta_{m}^{*}(0)$. Notice that above a particular value of $\beta_{m}$ the curves $\langle k_{m}\rangle/2n$ increase with $n$ in the panels in the top row of figure 7. This increase in mutual contacts characterizes a mixed phase where the two components share the same region in space (when the model transitions as $\beta_{m}$ increases beyond the critical point $\beta_{m}^{*}(0)\approx 0.31$ in figure 2 separating the SE- and M-phases).
@@
Figure 8: Determining the location of the critical point $\beta_{m}^{*}(0)$ from the locations of crossings between the $Var(k_{m})$ curves (denoted by $\beta_{m}^{*}(Var(k_{m})^{cr})$) and from the locations of crossing points of $d_{cm}$ curves (denoted by $\beta_{m}^{*}(d_{cm}^{cr})$) – see figure 7 panels (b1) and (b2), and panels (c1) and (c2). (a) The locations of the crossings of the variance per unit length curves $Var(k_{m})/2n$ as a function of $n=400$ with the corresponding curves for $m=48,148,200,300$, (b) As in the left panel but for the curves $d_{cm}$ against $\beta_{m}$. The three values of $\beta_{c}$ are $0.0,0.20$ and $0.28$ (see legend). In both panels data have been plotted against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ is the estimate of $\beta_{m}^{*}(0)$ (see reference [12]) for the $\beta_{c}=0$ case. These data appear to approach the critical point $\beta_{m}^{*}(0)$ as $n$ increases.
@@
Figure 9: (a1-a3) Plots of the average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, as a function of $\beta_{m}$ for $\beta_{c}=0.4,0.5$ and $0.75$. These values are well above the theta transition $\beta_{c}^{0}\approx 0.28$ (see the right panel of figure 4). Different symbols refer to different system sizes (see the legend). (b1-b3) Plots of the variance per unit length, $Var(k_{m})/2n$, as a function of $\beta_{m}$ for the same fixed values $\beta_{c}=0.4$, $0.5$ and $0.75$. (c1-c3) Corresponding $\beta_{m}$ dependence of the distance between the centre of mass of the two polygons, $d_{cm}$.
@@
Figure 10: Determining the locations of points on the critical curve $\beta_{m}^{*}(\beta_{c})$ by extrapolating the locations of crossings between the $Var(k_{m})$ curves (denoted by $\beta_{m}^{*}(Var(k_{m})^{cr})$) and extrapolating the locations of crossing points of $d_{cm}$ curves (denoted by $\beta_{m}^{*}(d_{cm}^{cr})$) – see figure 7 panels (b1) and (b2), and panels (c1) and (c2). and figure 9 panels (b1), (b2) and (b3), and panels (c1), (c2) and (c3). Left panel: The locations of crossings $\beta_{m}^{*}(Var(k_{m})^{cr})$ between the curve $Var(k_{m})/2n$ for $n=400$ (see the middle row of figure 9) and $Var(k_{m})/2n$ for $n\in\{48,100,148,200,300\}$ for values of $\beta_{c}<\beta_{c}^{0}$ (below the $\theta$-point; these are the empty symbols) with those above $\beta_{c}^{0}$ (solid symbols). The locations are plotted as a function of $1/\sqrt{n}$. Notice that the data below the $\theta$-point tends to line up and extrapolate as $n\to\infty$ to the marked solid vertical bar at $1/\sqrt{n}=0$ (this is the estimate for $\beta_{m}^{*}$ in reference [12] for the case $\beta_{c}=0$. In contrast, the data above the $\theta$-point extrapolate to larger values on the vertical axis. Right panel: This is similar to the left panel, but now for the curves $d_{cm}$ in the middle row of figure 9. The values of $\beta_{c}$ are reported in the legend.
@@
The onset of a thermodynamic phase transition between the SE- and M-phases is also suggested by the $\beta_{m}$ dependence of $Var(k_{m})/2n$ in the curves in the middle row of panels in figure 7. Notice the onset of peaks whose heights increase with $n$. Moreover, the curves cross one another in a narrow region of the mutual interaction parameter $\beta_{m}$. This is consistent with the theorem that, for $\beta_{c}<\infty$, and $\beta_{m}<\beta_{m}^{*}(0)$, $\lim_{n\to\infty}Var(k_{m})/2n=0$; see section 3 and theorem 1. That result shows the presence of an asymmetric phase transition when $\beta_{c}<\infty$ somewhat similar to the adsorption transition of a self-avoiding walk anchored to an impenetrable attractive wall [10]. In these cases it is known that the intersections of the curves $Var(k_{m})/2n$ plotted against $\beta_{m}$ provide a faithful finite-size estimate of the location of the phase boundary [15].
@@
A further indication of an SE-M phase boundary is seen in the behaviour of the distance between the centres of mass of the two components, $d_{cm}$, as a function of $\beta_{m}$ (these data are plotted in the bottom row panels in figure 7). In the SE phase, for $\beta_{c}=0$, the hard-core repulsion between vertices in the two polygons induces an entropic repulsive force between the polygons, segregating them when $\beta_{m}<\beta_{m}^{*}(0)$. Here the natural length scale is $n^{\nu}$, where $\nu$ is the metric exponent of the self-avoiding walk. This shows that $d_{cm}\sim n^{\nu}$ in the SE-phase. In the M-phase the two-component polygons share the same region as the attractive forces have overcome the entropic repulsion due to hard core nearest neighour interactions. In this case $d_{cm}$ should approach a constant, or grow slowly with $n$ at a rate slower than $O(n^{\nu})$. These expectations are supported by the data in the left-most panel in the bottom row of figure 7, and the corresponding data for $\beta_{c}=0.28$ seen in the other panel show consistent behaviour with $\beta_{c}=0$, also indicating that the phase boundary between the SE- and M-phases in figure 2 is a straight line.
@@
To determine the approximate location of the critical point $\beta_{m}^{*}(\beta_{c})$ the crossings between the curves $Var(k_{m})/2n$ and $d_{cm}$ were determined and plotted against $1/\sqrt{n}$ in figure 8 (assuming the mean field value $1/2$ for the finite size crossover exponent $\phi$ and plotting the intersections between the variance for $n=400$ and for $n=48,100,148,200$ and $300$). The data plausibly extrapolate to the value seen in reference [12], namely $\beta_{m}^{*}(0)\approx 0.31$ (the thick segment at $1/\sqrt{n}=0$ marks the estimate of the expanded-mixed phase transition $\beta_{m}^{*}(0)=0.31\pm 0.01$ reported in reference [12]). A similar extrapolation is reported for the location of the crossings of the curve $d_{cm}(n=400)$ with the corresponding curves obtained for $n=48,100,148,200$ and $300$ in the last row of panels in figure 8. The limiting estimate shown by the thick segment at $1/\sqrt{n}=0$ is again the expanded-mixed phase transition point $\beta_{m}^{*}(0)=0.31\pm 0.01$ reported in reference [12]. Similarly to the SE-SC phase boundary, we note that: (i) the $n\to\infty$ extrapolation of the points of both measures is very close to $\beta_{m}^{*}(0)$, and (ii) the SE-M phase transition appears to be independent of the strength $\beta_{c}$ of the self attraction between vertices in the polygons, at least when $0\leq\beta_{c}<\beta_{c}^{0}$.
@@
Figure 12: Determining the locations of points on the critical curve $\beta_{m}^{*}(\beta_{c})$ by extrapolating the locations of peaks in the $Var(k_{c})$ curves (denoted by $\beta_{c}^{*}(Var(k_{c})$) and extrapolating the locations of crossing points between $d_{cm}$ curves (denoted by $\beta_{c}^{*}(d_{cm}^{cr})$) – see figure 11 panels (b1), (b2) and (b3), and panels (e1), (e2) and (e3). (a) Comparison of the locations of the maxima of the variance curves $Var(k_{c})/2n$ for various values of $\beta_{m}$. These data plausibly extrapolate to $\beta_{c}^{0}$ for data below the M-SE transition (this corresponds to the empty symbols), but to larger values above the M-SE transition (solid symbols). (b) Similar to (a), but now for the separation between the centre-of-mass data $d_{cm}$ plotted for various values of $\beta_{m}$. The values of $\beta_{m}$ are reported explicitly in the legend. In both panels data have been plotted as a function of $1/\sqrt{n}$. The vertical bars at $1/\sqrt{n}=0$ refers to the estimates of $\beta_{c}^{0}$ provided in equation (1).